OpenStudy (anonymous):

PLEASE help! what is the correct simplification? (6^-2)^-3 I do not know how to start

OpenStudy (anonymous):

-2 x -3 = 6 6^6

OpenStudy (anonymous):

Could you help me with one more please?

OpenStudy (anonymous):

yes taking a power to a power you multiply the powers

OpenStudy (anonymous):

Thank you. 5^2/5^-6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry !! so this time it is division of two exponential terms: \[\frac{ 5^{2} }{ 5^{-6} }\] in case of division or multiplication, the exponents are subtracted or added to another. 2 -(-6) = 2+6 = 8 I believe answer is 5^8

OpenStudy (anonymous):

yes answer should be: \[5^{8}\]

OpenStudy (anonymous):

Thank you very much.